the plates of a parallel plate capacitor,5.0x10^-3m apart are maintained at a potential difference of 5.0x10^4 calculate the magnitude of the i. electric field intensity between the plates ii. force on the electron.
1
Expert's answer
2020-05-12T18:18:00-0400

For a parallel plate capacitor, "E=V\/d"

"\implies E=5*10^4/5*10^{-3}=10^7Vm^{-1}"

"Force =qE=1.6*10^{-19}*10^7=1.6*10^{-12} N"